find the y intercept, HA and the draw the shape
4) y=et - 2 5) y = 2e-* +1 6) y=-3.24 +1 7) y=-2* - 4 8) y=-e-* +3
Answer #1

Dus 4 X. е. 2 y = Solutiom! fo.x y - in tertept put x = 0 > y = e°-2 e-2=1-2 I y Z-1 > y. in the cep ti 10,-1) 2 =) Horizonta2 1 +0-1153-11solution : foru Ques 57 y = 26% +1 intere cept put xao > y = 2° +1 = 2+1=3 y - intere cepti (013) > Horizontal Asymptote y =y = 5.4 + 1 = 16.4 2 put oc. put oc=2 we ge t > y=a é y= 22?+ 1 = ax 0.1+1=121-2013 & -1 2 Y 6.41.611.2